### v1

# Formal Verification of Markov Processes with Learned Parameters

Abstract
We introduce the problem of formally verifying properties of Markov processes where the parameters are the output of machine learning models. Our formulation is general and solves a wide range of problems, including verifying properties of probabilistic programs that use machine learning, and subgroup analysis in healthcare modeling. We show that for a broad class of machine learning models, including linear models, tree-based models, and neural networks, verifying properties of Markov chains like reachability, hitting time, and total reward can be formulated as a bilinear program. We develop a decomposition and bound propagation scheme for solving the bilinear program and show through computational experiments that our method solves the problem to global optimality up to 100x faster than state-of-the-art solvers. We also release markovml, an open-source tool for building Markov processes, integrating pretrained machine learning models, and verifying their properties, available at https://github.com/mmaaz-git/markovml.
formal verification, Markov processes, machine learning, bilinear programming
1 Introduction
Markov processes are fundamental mathematical models used in diverse fields including computer science, operations research, engineering, and healthcare. In computer science, they can be used to model computer systems with probabilistic behavior, for example, hardware, communication protocols, or autonomous systems (Baier & Katoen, 2008). In engineering, they serve as fundamental models for analyzing machine degradation and failures (Rausand & Hoyland, 2003). In healthcare, they can model the progression of patients between health states, which can be used for cost-benefit analysis of drugs and medical technologies (Sonnenberg & Beck, 1993). We refer the reader to Stewart (2021) for a comprehensive review of their applications.
Increasingly, machine learning (ML) models are being integrated into probabilistic systems either as components of the systems themselves, e.g., integrating ML models into autonomous systems (Baier & Katoen, 2008), or to enrich the Markov models by incorporating heterogeneous parameters from observed data. For example, an ML model can predict failure rates based on machine-specific conditions and these predictions can be incorporated into a Markov chain to determine expected lifetime. In healthcare, transition probabilities can be patient specific, depending on demographics or specific clinical features (Mertens et al., 2022). Effectively, this leads to Markov processes where the parameters are not fixed, but instead are represented by learned functions.
There are few approaches to rigorously analyzing Markov processes with embedded ML models. In the healthcare literature, such Markov models are most often analyzed using simulation (Krijkamp et al., 2018). While this allows for subgroup analysis, e.g., performing a cost-benefit analysis for men over the age of 60, simulation cannot provide formal guarantees about outcomes. This limitation is particularly important in high-stakes applications like healthcare decision-making or safety-critical systems.
In this paper, we develop a new approach to analyzing Markov processes with embedded ML models. Our key insight is that techniques developed for formal verification of ML models can be extended to analyze Markov processes with embedded ML models. This connection allows us to move beyond simulation to provide formal guarantees for properties of Markov processes by solving a bilinear program. While our primary motivation comes from healthcare applications, our framework is general and applies to any domain where ML models are used as inputs to Markov processes. This allows us to answer questions like:
  • For a probabilistic program with a given bound on the input, what is the maximum probability of reaching a failure state?
  • For machines with certain properties, is the probability of failure below 0.01%?
  • For patients within a certain subgroup, is the expected cost of a drug below the government’s threshold?
To summarize, our main contributions are:
  1. 1.
    We develop a general framework for formally verifying properties of Markov processes with parameters determined by ML models.
  2. 2.
    We show that, for a broad class of ML models including linear models, tree-based models, and ReLU-based neural networks, we can formulate common properties of Markov processes as a bilinear program.
  3. 3.
    We develop a decomposition and bound propagation scheme for solving the resulting bilinear program and show that we can solve the problem to global optimality several orders of magnitude faster than state-of-the-art solvers.
2 Related Work
2.1 Formal Verification of Markov Processes
Probabilistic model checking is a set of techniques used in computer science to formally verify the properties of probabilistic systems, with foundations in Markov chain theory (Hansson & Jonsson, 1994; Baier & Katoen, 2008). Tools like PRISM (Kwiatkowska et al., 2002), MRMC (Katoen et al., 2005), $\text{E}\vdash\text{MC}^{2}$ (Hermanns et al., 2000), VESTA (Sen et al., 2005), YMER (Younes, 2005), and APMC (Lassaigne & Peyronnet, 2002) allow users to define a Markov process with fixed parameters and verify whether specific properties hold using numerical methods. Some tools, such as PRISM, also support parametric model checking (Chen et al., 2013), where the parameters of the Markov process are represented as rational functions of external variables. In this light, our work broadens the scope of parametric model checking to situations when the parameters are determined by ML models. Similar to current model checking tools, we developed an expressive domain-specific language that allows users to specify and verify properties of Markov processes with learned parameters.
2.2 Markov Processes with Uncertain Parameters
Markov processes with variable parameters are well-studied in the mathematics and operations research literature. There are generalizations of classical Markov theory results, like the Perron-Frobenius theorem, bounds on hitting times, and bounds on stationary distributions to when the transition matrix lies in a credal set (Blanc & Den Hertog, 2008; De Cooman et al., 2014). Marbach & Tsitsiklis (2001, 2003) study Markov reward processes where the probabilities are a differentiable function of a parameter. There is also a large body of work on Markov decision processes with uncertain parameters, well known in the reinforcement learning literature (El Ghaoui & Nilim, 2005; Iyengar, 2005; Goyal & Grand-Clement, 2023; Grand-Clément & Petrik, 2024). Goh et al. (2018), also inspired by healthcare modeling, develop a value iteration algorithm to compute the optimum of the total reward of a Markov process when the transition matrix lies in an uncertainty set. Chan & Maaz (2024) study the “inverse” problem of finding an algebraic description of all parameters for which the total reward obeys a certain threshold. In contrast to these papers, our paper studies uncertain parameters which are given by ML models.
2.3 Formal Verification of ML Models
Formal verification of ML models has emerged as an important research area in recent years, particularly for safety-critical applications. Here, the properties to verify are, e.g., robustness to input perturbations, or bounds on the output. Most of this work has been driven by interest in verifying properties of neural networks, as they are complex, widely deployed in critical applications like aircraft control (Owen et al., 2019), yet can be sensitive to small perturbations in input (Szegedy, 2013). Neural networks with ReLU activations can be easily encoded into SMT (satisfiability modulo theories) or into a mixed-integer linear program. A key breakthrough was the development of Reluplex (Katz et al., 2017), a modified simplex algorithm for verifying properties of neural networks. Modern state-of-the-art approaches, like $\alpha,\beta$-CROWN, integrate various methods like bound propagation, branch-and-bound, and GPU-optimized implementations (Wang et al., 2021; Zhang et al., 2022b, a; Kotha et al., 2023). There is increasing interest in the optimization literature on optimal formulations of ML models, particularly for neural networks (Anderson et al., 2020; Tjandraatmadja et al., 2020; Kronqvist et al., 2021; Anh-Nguyen & Huchette, 2022). Our paper leverages these formulations to embed ML models into the overall optimization problem.
2.4 Subgroup Analysis in Healthcare
Our work is most directly inspired by the problem of subgroup analysis in healthcare cost-benefit analysis. Markov “microsimulation” models are a popular tool for analyzing the trajectories of patients over time, e.g., the progression of cancer or other diseases, which allow for heterogeneous parameters (Krijkamp et al., 2018). These allow for the computation of important metrics in healthcare economics like the incremental cost-effectiveness ratio (ICER) and the net monetary benefit (NMB) (Sonnenberg & Beck, 1993). In a scoping review, Mertens et al. (2022) find that estimation of individual disease risk – in Markov theory, computing the transition probabilities – is often computed using so-called risk scores from the medical literature (e.g., (Wilde et al., 2019; Lee et al., 2020; Breeze et al., 2017)). These risk score formulas are most often computed with logistic regression applied to observational data, though other models are beginning to be used. Most current work in this domain uses Monte Carlo simulation (Krijkamp et al., 2018). In contrast, our work allows provides an exact solution to subgroup analysis by solving an optimization problem to global optimality. Our framework also allows for the analysis of a broad class of ML models, beyond just logistic regression, thus ensuring applicability as healthcare analytics continues to adopt more sophisticated machine learning approaches.
2.5 Bilinear Programming
As we will show, our problem can be formulated as a bilinear program. Bilinear programs are a type of non-convex NP-hard optimization problem where the objective and/or the constraints are bilinear functions (Horst & Pardalos, 2013). In theory, it is possible solve them to global optimality with branch-and-bound methods (Horst & Pardalos, 2013), among others (Gallo & Ülkücü, 1977). The McCormick relaxation (McCormick, 1976) is a standard tool for bilinear programs which forms a convex relaxation, helping the solver to obtain good bounds on the solution. Indeed, this and other methods are already implemented in modern optimization software like Gurobi (Gurobi Optimization, LLC, 2024). Under additional assumptions on the structure of the bilinear program, it is possible to form tighter relaxations than McCormick’s (Dey et al., 2019), or even solve the optimization problem using mixed-integer linear programming (Horst & Pardalos, 2013). Unfortunately, our problem does not satisfy these structural requirements and thus requires a new approach.
3 Problem Formulation
3.1 Notation
Vectors are lowercase bold, e.g., $\boldsymbol{\mathbf{x}}$, with $i$th entry $x_{i}$, and matrices by uppercase bold letters, e.g., $\boldsymbol{\mathbf{M}}$ with $(i,j)$-th entry $M_{ij}$. The identity matrix is denoted by $\boldsymbol{\mathbf{I}}$, the vector of all ones by $\boldsymbol{\mathbf{1}}$, with dimensions inferred from context. The set of integers from $1$ to $n$ is denoted by $[n]$.
3.2 Preliminaries
A (discrete-time, finite-state) Markov chain with $n$ states, indexed by $[n]$, is defined by a row-stochastic transition matrix $\boldsymbol{\mathbf{P}}\in\mathbb{R}^{n\times n}$, where $P_{ij}$ is the probability of transitioning from state $i$ to state $j$, and a stochastic initial distribution vector $\boldsymbol{\mathbf{\pi}}\in\mathbb{R}^{n}$, where $\pi_{i}$ is the probability of starting in state $i$. Furthermore, if we assign rewards to each state, we call this a Markov reward process. A Markov reward process has a reward vector $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{n}$, where $r_{i}$ is the reward for being in state $i$ for one period. A state is absorbing if it cannot transition to any other state, and transient otherwise.
Below, we recount three of the key properties commonly computed in practice (Puterman, 2014).
Definition 3.1 (Reachability).
The probability of eventually reaching a set of states $S\subseteq[n]$, from a set $T\subseteq[n]$ of transient states not in $S$, is given by $\boldsymbol{\mathbf{\tilde{\pi}}}^{\top}(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$, where $\boldsymbol{\mathbf{Q}}$ is the transition matrix restricted to $T$, $\boldsymbol{\mathbf{R}}$ is the transition matrix from $T$ to $S$, and $\boldsymbol{\mathbf{\tilde{\pi}}}$ is the initial distribution over $T$.
Definition 3.2 (Expected hitting time).
The expected number of steps to eventually reach a set of states $S\subseteq[n]$, from a set $T\subseteq[n]$ of transient states not in $S$, assuming that the chain will reach $S$ from $T$ with probability 1, is given by $\boldsymbol{\mathbf{\tilde{\pi}}}^{\top}(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{1}}$, where $\boldsymbol{\mathbf{Q}}$ is the transition matrix restricted to $T$, and $\boldsymbol{\mathbf{\tilde{\pi}}}$ is the initial distribution over $T$.
Definition 3.3 (Total infinite-horizon discounted reward).
The total infinite-horizon discounted reward, with a discount factor $\lambda\in(0,1)$ is given by $\sum_{t=0}^{\infty}\lambda^{t}\boldsymbol{\mathbf{\pi}}^{\top}\boldsymbol{\mathbf{P}}^{t}\boldsymbol{\mathbf{r}}=\boldsymbol{\mathbf{\pi}}^{\top}\left(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}}\right)^{-1}\boldsymbol{\mathbf{r}}$.
For each of the above quantities, we can restrict the analysis to a single state, e.g., for reachability starting from a state $i$, we simply set $\tilde{\pi}_{i}=1$ and for $j\neq i$, set $\tilde{\pi}_{i}=0$.
These three quantities enable rich analysis of Markov processes and the systems they model. Reachability can be used to verify the probability of failure in a system. Expected hitting time can be used to compute the expected time to failure, or life expectancy of a person. Total reward can be used to compute the resource consumption of a system or total cost of a drug.
We wish to verify these properties, namely finding their maximum or minimum. If the Markov process’ parameters are fixed, as is common in model checking, these quantities can be computed exactly by solving a linear system. However, in our case, the parameters are given by ML models, and so we will formulate an optimization problem to derive bounds on these quantities.
As the three quantities have similar formulations (e.g., hitting time is essentially the same as the total reward if the $\boldsymbol{\mathbf{r}}$ is replaced by $\boldsymbol{\mathbf{1}}$ and suitable adjustments for the initial distribution and transition matrix), we will proceed in the rest of the paper by studying the total reward. It will be easy to modify results for the other two quantities.
3.3 Embedding ML Models
We now consider $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ as functions of a feature vector $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{m}$, where $m$ is the number of features.
We encode the relationship from $\boldsymbol{\mathbf{x}}$ to $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ via a set of functions, $f_{1},f_{2},\ldots,f_{k_{f}}$, where $k_{f}$ is the number of functions. Each function $f_{i}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{\ell_{i}}$ for $i=1,2,\ldots,k_{f}$ takes a feature vector and outputs a vector $\boldsymbol{\mathbf{\theta_{i}}}$ (e.g., a classifier may output class probabilities for three classes). We concatenate these vectors to form the output vector $\boldsymbol{\mathbf{\theta}}\in\mathbb{R}^{\ell}$, where $\ell=\sum_{i=1}^{k_{f}}\ell_{i}$.
Our key assumption on the functions will be that they are mixed-integer linear representable (MILP-representable), meaning that the relationship between the inputs and outputs can be expressed using linear constraints and binary variables. This is a broad class of functions that includes, e.g., piecewise linear functions (including simple “if-then” rules), linear and logistic regression, tree-based models, and neural networks with ReLU activations [^1] A caveat is that in the case of logistic regression and neural network classifiers, there is a non-linearity introduced by the softmax or logistic function. Typically in the formal verification literature, to check, e.g., robustness, it suffices to check the logits, not the actual probabilities. However, in our problem, we need the probabilities as they are inputs into our Markov process. Practically, we will handle this issue using the adaptive piecewise linear approximations available in modern optimization solvers..
Next, the output vector $\boldsymbol{\mathbf{\theta}}$ is linked to $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}}$ and $\boldsymbol{\mathbf{r}}$ through affine equalities: $\boldsymbol{\mathbf{\pi}}=\boldsymbol{\mathbf{A_{\pi}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\pi}}}$, $\operatorname{vec}(\boldsymbol{\mathbf{P}})=\boldsymbol{\mathbf{A_{P}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{P}}}$, and $\boldsymbol{\mathbf{r}}=\boldsymbol{\mathbf{A_{r}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{r}}}$, for fixed $\boldsymbol{\mathbf{A_{\pi}}}\in\mathbb{R}^{n\times\ell},\boldsymbol{\mathbf{A_{P}}}\in\mathbb{R}^{n^{2}\times\ell},\boldsymbol{\mathbf{A_{r}}}\in\mathbb{R}^{n\times\ell}$ and $\boldsymbol{\mathbf{b_{\pi}}}\in\mathbb{R}^{n},\boldsymbol{\mathbf{b_{P}}}\in\mathbb{R}^{n^{2}},\boldsymbol{\mathbf{b_{r}}}\in\mathbb{R}^{n}$. Note the sizes of the matrices mean that every element of $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ must be bound with an affine equality to $\boldsymbol{\mathbf{\theta}}$. In other words, if we fix $\boldsymbol{\mathbf{\theta}}$, we can compute $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ exactly. The $\operatorname{vec}$ operator vectorizes a matrix, i.e., by concatenating rows into one column vector, so that $\operatorname{vec}(\boldsymbol{\mathbf{P}})\in\mathbb{R}^{n^{2}}$.
The affine equalities allow for common formulations like one of the parameters equaling an element of $\boldsymbol{\mathbf{\theta}}$, e.g., $\pi_{1}=\theta_{1}$. Another common situation is that we have a function that outputs the probability of transitioning to a specific state (say, a death state in a healthcare model), and the remaining probability mass is, e.g., distributed uniformly over the remaining states: $\pi_{n}=\theta_{1}$ and $\pi_{i}=(1-\sum_{j\neq i}\theta_{j})/(n-1)$ for $i\neq n$. Of course, the affine equalities also allow for parameters to be fixed as constants.
Next, we may have linear inequalities on the parameters: $\boldsymbol{\mathbf{C_{\pi}}}\boldsymbol{\mathbf{\pi}}\leq\boldsymbol{\mathbf{d_{\pi}}}$, $\boldsymbol{\mathbf{C_{P}}}\operatorname{vec}(\boldsymbol{\mathbf{P}})\leq\boldsymbol{\mathbf{d_{P}}}$, and $\boldsymbol{\mathbf{C_{r}}}\boldsymbol{\mathbf{r}}\leq\boldsymbol{\mathbf{d_{r}}}$, for fixed $\boldsymbol{\mathbf{C_{\pi}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{\pi}}}\times n},\boldsymbol{\mathbf{C_{P}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{P}}}\times n^{2}},\boldsymbol{\mathbf{C_{r}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{r}}}\times n}$ and $\boldsymbol{\mathbf{d_{\pi}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{\pi}}}},\boldsymbol{\mathbf{d_{P}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{P}}}},\boldsymbol{\mathbf{d_{r}}}\in\mathbb{R}^{k_{\boldsymbol{\mathbf{r}}}}$, where $k_{\boldsymbol{\mathbf{\pi}}},k_{\boldsymbol{\mathbf{P}}},k_{\boldsymbol{\mathbf{r}}}$ are the number of linear inequalities on $\boldsymbol{\mathbf{\pi}}$, $\boldsymbol{\mathbf{P}}$, and $\boldsymbol{\mathbf{r}}$, respectively.
Linear inequalities are an appropriately general assumption as they arise naturally from fixing a monotonic ordering on the rewards or the probabilities. In reliability engineering and healthcare, a common assumption is that of increasing failure rate, which implies a certain stochastic ordering on the transition probabilities (Barlow & Proschan, 1996; Alagoz et al., 2007). Effectively, this introduces a series of linear inequalities on the transition probabilities. See Figure 1 for a diagram of the full pipeline.
Lastly, we assume that the feature vector $\boldsymbol{\mathbf{x}}$ lies in a set $\mathcal{X}\subseteq\mathbb{R}^{m}$. We assume that the set $\mathcal{X}$ is mixed-integer-linear representable (MILP-representable), i.e., it can be written as a finite union of polyhedra (Jeroslow & Lowe, 1984). This is a common assumption in the optimization and ML literature, and is satisfied by a wide range of sets, including box constraints, polytopes, and discrete sets. In practice, this may look like, e.g., that we restrict $\boldsymbol{\mathbf{x}}$ to represent the set of patients that are either under 18 or over 65, as this is a union of two intervals.
Figure 1: Example of our pipeline. A feature vector $\boldsymbol{\mathbf{x}}$ is passed through different functions, here a linear regression and a neural network, to obtain the output vector $\boldsymbol{\mathbf{\theta}}$, which then determines the parameters of the Markov process through affine equalities.
With these assumptions, we can now formulate the following optimization problem for the total infinite-horizon discounted reward:
[…]
$\quad\boldsymbol{\mathbf{v}}=\lambda\boldsymbol{\mathbf{P}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{r}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\pi}}=\boldsymbol{\mathbf{A_{\pi}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\pi}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{P}})=\boldsymbol{\mathbf{A_{P}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{P}}},$
$\quad\boldsymbol{\mathbf{r}}=\boldsymbol{\mathbf{A_{r}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{r}}},$
$\quad\boldsymbol{\mathbf{C_{\pi}}}\boldsymbol{\mathbf{\pi}}\leq\boldsymbol{\mathbf{d_{\pi}}},$
$\quad\boldsymbol{\mathbf{C_{P}}}\operatorname{vec}(\boldsymbol{\mathbf{P}})\leq\boldsymbol{\mathbf{d_{P}}},$
$\quad\boldsymbol{\mathbf{C_{r}}}\boldsymbol{\mathbf{r}}\leq\boldsymbol{\mathbf{d_{r}}},$
$\quad\sum_{j=1}^{n}P_{ij}=1\quad\forall i\in[n],$
$\quad\sum_{i=1}^{n}\pi_{i}=1,$
$\quad 0\leq P_{ij}\leq 1\quad\forall i,j\in[n]\times[n],$
$\quad 0\leq\pi_{i}\leq 1\quad\forall i\in[n],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
Above, $\boldsymbol{\mathbf{v}}$ is a new vector we have introduced, which in Markov process theory would be referred to as the value function, and often arises in reinforcement learning, which equals $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}\boldsymbol{\mathbf{r}}$. We can easily convert this into a feasibility problem, to find a feasible $\boldsymbol{\mathbf{x}}$, by replacing the objective function with an inequality, $W_{\min}\leq\boldsymbol{\mathbf{\pi}}^{\top}\boldsymbol{\mathbf{v\leq W_{\max}}}$. With slight modifications, we can also formulate reachability and hitting time: the full formulations are in Appendix A.
4 Solving the Optimization Problem
Under the assumptions that the functions $f_{1},f_{2},\ldots,f_{k_{f}}$ are MILP-representable and $\mathcal{X}$ is MILP-representable, the optimization problem above is a mixed-integer bilinear program, as we have a bilinear objective and a bilinear constraint. We will leverage the observation that strong bounds on variables are crucial for solving bilinear programs, and will derive such bounds by decomposing our problem and using various results from Markov theory.
Our strategy is as follows: obtain bounds on the elements of $\boldsymbol{\mathbf{\theta}}$ by solving a series of smaller MILPs, propagate these bounds to the parameters via the affine equalities, obtain bounds on $\boldsymbol{\mathbf{v}}$ using linear algebra arguments, and then tighten the $\boldsymbol{\mathbf{v}}$ bounds using interval matrix analysis. All proofs are in Appendix B
4.1 Bounds on $\boldsymbol{\mathbf{\theta}}$
For each $\theta_{i}$, we minimize and maximize it by solving two smaller MILPs, over the set $\mathcal{X}$. Namely, for each $i\in[k_{f}]$, for each $j\in[\ell_{i}]$, solve for $\mathop{\text{min}/\text{max}}_{\boldsymbol{\mathbf{x}}\in\mathcal{X}}\theta_{i,j}\;\text{s.t.}\;\boldsymbol{\mathbf{\theta_{i}}}=[\theta_{i,1},\ldots,\theta_{i,\ell_{i}}]=f_{i}(\boldsymbol{\mathbf{x}})$, which are MILPs as $\mathcal{X}$ and $f_{1},\ldots f_{k_{f}}$ are MILP-representable.
4.2 Bounds on $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$
We now propagate the bounds on $\boldsymbol{\mathbf{\theta}}$ to the parameters $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ via the affine equalities.
Lemma 4.1.
Let $\boldsymbol{\mathbf{\theta}}\in\mathbb{R}^{\ell}$ be a vector where each component satisfies the bounds $\theta_{i}^{\min}\leq\theta_{i}\leq\theta_{i}^{\max}$ for all $i=1,2,\ldots,\ell$. Consider an affine transformation defined by $\boldsymbol{\mathbf{y}}=\boldsymbol{\mathbf{A}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b}}$, where $\boldsymbol{\mathbf{y}}\in\mathbb{R}^{k}$, $\boldsymbol{\mathbf{A}}\in\mathbb{R}^{k\times\ell}$, and $\boldsymbol{\mathbf{b}}\in\mathbb{R}^{k}$. Then, each component $y_{i}$ of $\boldsymbol{\mathbf{y}}$ satisfies $b_{i}+\sum_{j=1}^{\ell}A_{ij}\cdot\left(\theta_{j}^{\min}\mathbbm{1}_{A_{ij}\geq 0}+\theta_{j}^{\max}\mathbbm{1}_{A_{ij}<0}\right)\leq y_{i}\leq b_{i}+\sum_{j=1}^{\ell}A_{ij}\cdot\left(\theta_{j}^{\max}\mathbbm{1}_{A_{ij}\geq 0}+\theta_{j}^{\min}\mathbbm{1}_{A_{ij}<0}\right)$.
4.3 Bounds on $\boldsymbol{\mathbf{v}}$
Next, we derive bounds on $\boldsymbol{\mathbf{v}}$ using well-established facts from Markov theory. First, as $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}\boldsymbol{\mathbf{r}}$, we analyze bounds on $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}$.
Lemma 4.2.
Let $\boldsymbol{\mathbf{P}}$ be row-stochastic and $\lambda\in(0,1)$. Then, each element in $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}$ is in the interval $[0,1/(1-\lambda)]$ and the row sums are all equal to $1/(1-\lambda)$.
Now we can propagate the above bounds to $\boldsymbol{\mathbf{v}}$.
Lemma 4.3.
Let $r_{i}^{\min}\leq r_{i}\leq r_{i}^{\max}$. Then, each element in the vector $\boldsymbol{\mathbf{v}}$ satisfies $\gamma\min_{j}r_{j}^{\min}\leq v_{i}\leq\gamma\max_{j}r_{j}^{\max}$, where $\gamma=\frac{1}{1-\lambda}$.
Lemma 4.3 provides valid bounds on $\boldsymbol{\mathbf{v}}$. However, these bounds are not tight, as Lemma 4.2 applies to any row-stochastic matrix, and does not take into account the specific element-wise bounds on $\boldsymbol{\mathbf{P}}$ that we derived from Lemma 4.1. In order to derive tighter bounds on $\boldsymbol{\mathbf{v}}$, we need to incorporate the specific element-wise bounds on $\boldsymbol{\mathbf{P}}$. This is difficult because of the matrix inverse. In the following section, we show how to solve this problem with interval matrix analysis and hence derive tighter bounds on $\boldsymbol{\mathbf{v}}$.
4.4 Tightening bounds on $\boldsymbol{\mathbf{v}}$
Interval matrix analysis studies a generalization of matrices (or vectors), where each element is an interval instead of a real number. Equivalently, it can be viewed as studying the set of all matrices that lie within given intervals. There are easy generalizations of arithmetic of real numbers to intervals, which can be used to generalize matrix arithmetic to interval matrices. However, inverting an interval matrix is difficult due to the inherent nonlinearity of the inverse operation. Consider the equivalent problem of solving a linear system with interval matrices. The solution set is not rectangular, so cannot be described as an interval matrix. Hence, we wish to find the interval matrix of smallest radius that contains the solution set, known as the hull (Neumaier, 1984).
Due to the interval extensions of the basic arithmetic operations, it is possible to generalize the Gauss-Seidel method, a common iterative procedures for solving linear systems, to interval matrices. Beginning with an initial enclosure of the solution set, it iteratively refines the intervals and returns a smaller enclosure of the solution set. We provide a full description of the Gauss-Seidel method in Appendix C. The interval Gauss-Seidel method is optimal in the sense that it provides the smallest enclosure of the solution set out of a broad class of algorithms (Neumaier, 1984).
Hence, we can apply the Gauss-Seidel method to tighten the bounds on $\boldsymbol{\mathbf{v}}$, as $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{r}}$ is a linear system. We use the bounds on $\boldsymbol{\mathbf{P}}$ and $\boldsymbol{\mathbf{r}}$ from Lemma 4.1 to form the interval matrices, and for our initial guess, we can use the bounds derived from Lemma 4.3222Typically, one has to be careful with Gauss-Seidel because it involves a division, and division by an interval containing zero is not defined. However, this situation will never occur in our problem, as we discuss in Appendix C, rendering it particularly amenable to Gauss-Seidel..
Still, Gauss-Seidel is not guaranteed to obtain the hull itself. A well-known sufficient condition for Gauss-Seidel to obtain the hull is that the matrix in the linear system is an interval M-matrix, a generalization of the notion of an M-matrix (Berman & Plemmons, 1994) to interval matrices. Suppose a matrix $\boldsymbol{\mathbf{M}}$ is element-wise bounded by $\boldsymbol{\mathbf{M^{\max}}}$ and $\boldsymbol{\mathbf{M^{\min}}}$. It is an interval M-matrix if and only if $\boldsymbol{\mathbf{M^{\min}}}$ is a (non-singular) real M-matrix in the usual sense and $\boldsymbol{\mathbf{M^{\max}}}$ has no positive off-diagonal elements (Neumaier, 1984). Hence we can prove a necessary and sufficient condition for $\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}}$ to be an interval M-matrix.
Theorem 4.4.
Let $\boldsymbol{\mathbf{P}}$ be a row-stochastic matrix bounded element-wise by $\boldsymbol{\mathbf{P^{\min}}}\leq\boldsymbol{\mathbf{P}}\leq\boldsymbol{\mathbf{P^{\max}}}$. Then, $\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}}$ is an interval M-matrix if and only if $\rho(\boldsymbol{\mathbf{P}}^{\max})\leq\frac{1}{\lambda}$, where $\rho$ is the spectral radius.
Notably, we only need check the spectral radius of the matrix formed by the upper bounds on $\boldsymbol{\mathbf{P}}$. There are various conditions that can be used to obtain the necessary bound on the spectral radius. Since $\lambda<1$, we have $11/\lambda>$ 1, so if we can show that the spectral radius of the upper bound matrix is less than $1$, then this automatically implies the condition in Theorem 4.4, and hence that the Gauss-Seidel method will obtain the hull, i.e., the tightest possible rectangular bounds on $\boldsymbol{\mathbf{v}}$. This may occur if, for example, the row sums of the upper bounds are $\leq 1$.
Example 4.5.
Let $\lambda=0.97$, and suppose we have the following bounds on $\boldsymbol{\mathbf{P}}$ and $\boldsymbol{\mathbf{r}}$:
$\boldsymbol{\mathbf{P}}:\begin{bmatrix}[0.5,0.6]&[0.2,0.5]\\ [0.1,0.4]&[0.5,0.6]\end{bmatrix},\quad\boldsymbol{\mathbf{r}}:\begin{bmatrix}[0,100]\\ [0,100]\end{bmatrix}$
The bounds on $\boldsymbol{\mathbf{v}}$ are:
[…]
Note that the spectral radius of the upper bound matrix of $\boldsymbol{\mathbf{P}}$ is $1.05$, which is greater than $1.031/\lambda=1/0.97=$ 1.03. So, Gauss-Seidel will not necessarily obtain the hull, as it is not an interval M-matrix, by Theorem 4.4.
If we slightly tighten the upper bounds on $\boldsymbol{\mathbf{P}}$ by changing the upper bound on $P_{2,1}$ to $0.3$, then the bounds on $\boldsymbol{\mathbf{v}}$ are:
[…]
Indeed, now the spectral radius of the upper bound matrix of $\boldsymbol{\mathbf{P}}$ is $0.99$, which is less than $1.031/\lambda=$ 1.03, so by Theorem 4.4, Gauss-Seidel returns the hull. Note the tighter bounds than the prior case.
4.5 Solving the final optimization problem
At last, we can assert the bounds found above for $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}},\boldsymbol{\mathbf{v}}$ as constraints and solve the full bilinear program with any out-of-the-box solver. We summarize our procedure in Algorithm 1.
Algorithm 1 Decomposition and bound propagation scheme to optimize the infinite-horizon total reward
  Input: Instance of the infinite-horizon total reward optimization problem
  for $i=1$ to $k_{f}$ do
     for $j=1$ to $\ell_{i}$ do
        Minimize and maximize $\theta_{i,j}$ over $\mathcal{X}$
     end for
  end for
  Propagate the bounds on $\boldsymbol{\mathbf{\theta}}$ to $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$ using Lemma 4.1
  Obtain initial bounds on $\boldsymbol{\mathbf{v}}$ using Lemma 4.3
  Tighten the bounds on $\boldsymbol{\mathbf{v}}$ using Gauss-Seidel
  Optimize $\boldsymbol{\mathbf{\pi}}^{\top}\boldsymbol{\mathbf{v}}$ with an out-of-the-box solver, asserting the obtained bounds on $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}},\boldsymbol{\mathbf{v}}$ as constraints
Our algorithm works with any standard bilinear or MILP solver, avoiding the need for specialized branch-and-bound schemes or custom callbacks that require expert implementation. The bound propagation and Gauss-Seidel are also straightforward to implement.
4.6 Extensions: reachability and hitting time
Our results also extend to reachability and hitting time, if we take care of minor technical details arising from the invertibility of $\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}}$. See Appendix D for full details.
5 Implementation
We developed the Python package markovml to specify Markov chains or reward processes with embedded pretrained machine learning models. Our domain-specific language lets users:
  1. 1.
    Instantiate a Markov process
  2. 2.
    Add pretrained ML models from sklearn (Pedregosa et al., 2011) and PyTorch (Paszke et al., 2019)
  3. 3.
    Link model outputs to Markov parameters with affine equalities
  4. 4.
    Include extra linear inequalities
  5. 5.
    Specify the feature set with MILP constraints
  6. 6.
    Optimize reachability, hitting time, or total reward
Built on gurobipy (Gurobi Optimization, LLC, 2024), our package supports a variety of regression and classification models, including linear, tree-based, and neural networks. For some models, the MILP formulation is provided by gurobi-machinelearning (Gurobi Optimization, 2024), while for some models (e.g., softmax for classifiers), we implemented the MILP formulation ourselves. We also allow the user to build “if-then” rules in a natural syntax. Our package fully implements Algorithm 1. See Appendix E for more details.
6 Numerical experiments
We compare our method to solving the optimization problem directly with an out-of-the-box solver. Our experiments will analyze the following drivers of complexity of the bilinear program: the number of states, the number of ML models, and the model complexity, notably for trees and neural networks.
6.1 Setup
For all experiments, we fix the number of features $m=5$, the feature set $\mathcal{X}=[-1,1]^{5}$, and the discount factor $\lambda=0.97$. We train ML models on randomly generated data ($10,000$ random points). For regression models, we draw $\boldsymbol{\mathbf{X}}\sim\mathcal{N}(0,1)^{10000\times 5}$, $\boldsymbol{\mathbf{\beta}}\sim\mathcal{N}(0,1)^{5}$, and then generate data points as $\boldsymbol{\mathbf{X}}\boldsymbol{\mathbf{\beta}}+\boldsymbol{\mathbf{\epsilon}}$, where $\boldsymbol{\mathbf{\epsilon}}\sim\mathcal{N}(0,1)^{10000}$. For classification models, we use the same linear form but draw data points from a Bernoulli distribution with probabilities $\sigma(\boldsymbol{\mathbf{X}}\boldsymbol{\mathbf{\beta}})$, where $\sigma(\cdot)$ is the logistic function.
For rewards, we train a regression model whose output is set to $r_{1}$. For $i=2,\ldots,n$, we set $r_{i}$ to be output divided by $i$. For probabilities, we train binary classifiers for $\boldsymbol{\mathbf{\pi}}$ and rows of $\boldsymbol{\mathbf{P}}$. For $\boldsymbol{\mathbf{\pi}}$, we set $\pi_{1}$ equal to the classifier output, the remaining probability is assigned to $\pi_{2}$, and the remaining are set to zero. For rows of $\boldsymbol{\mathbf{P}}$, the classifier output is the probability of remaining in the same state, with the remaining probability assigned to transitioning to the next state, and the rest of the row is 00. We make the last state absorbing, i.e., transitions to itself with probability $1$.
For each configuration, we run 10 instances, each with random data, train the models on the random data, and solve the optimization problem directly with an out-of-the-box solver and with our method.
We perform the following experiments:
  1. 1.
    State Space: Vary $n\in\{5,10,20,50,100,200\}$ using three models: a linear regression for $\boldsymbol{\mathbf{r}}$ and two logistic regressions for $\boldsymbol{\mathbf{\pi}}$ and the first row of $\boldsymbol{\mathbf{P}}$ (the remaining rows are uniform).
  2. 2.
    Model Count: Fix $n=20$ with linear regression for $\boldsymbol{\mathbf{r}}$ and logistic regression for all probability models. Models for $\boldsymbol{\mathbf{r}}$ and $\boldsymbol{\mathbf{\pi}}$ are always used, while the number of modeled rows in $\boldsymbol{\mathbf{P}}$ varies from $1$ to $19$, so that the total number of models runs from $3$ to $21$.
  3. 3.
    Decision Tree Depth: Fix $n=20$ using three decision tree models (for $\boldsymbol{\mathbf{r}}$, $\boldsymbol{\mathbf{\pi}}$, and the first row of $\boldsymbol{\mathbf{P}}$) with depths in $\{2,3,4,6,8\}$. Remaining rows of $\boldsymbol{\mathbf{P}}$ are uniform.
  4. 4.
    Neural Network Architecture: Fix $n=20$ using three ReLU multilayer perceptrons, for $\boldsymbol{\mathbf{r}}$, $\boldsymbol{\mathbf{\pi}}$, and the first row of $\boldsymbol{\mathbf{P}}$. We vary hidden layers in $\{1,2\}$ and neurons per layer in $\{5,10,15,20\}$, with remaining rows of $\boldsymbol{\mathbf{P}}$ set uniformly.
Figure 2: Runtimes of our method versus directly solving. Each panel shows results from experiments 1-4 (left to right). Points represent mean values with error bars indicating ±1 standard deviation. Statistical significance from paired t-tests is denoted above: * ($p<0.05$), † ($p<0.01$), and § ($p<0.001$).
Figure 3: Proportion (%) of instances categorized as optimal, suboptimal, timed out, or other. Each panel shows experiments 1-4 (left to right). Our method is the left bar and direct method is the right bar. Statistical significance from $\chi^{2}$ tests is denoted above: * ($p<0.05$), † ($p<0.01$), and § ($p<0.001$).
We record runtime, objective value, and optimizer status. For our method, the runtime includes solving the smaller MILPs, bound computations, the interval Gauss-Seidel phase, and the final optimization. All experiments are run on an Intel Core i7 with 6 cores using Gurobi Optimizer 12.0.0 (Gurobi Optimization, LLC, 2024) (1200 second limit per problem, and presolve disabled throughout). All code and instructions on how to reproduce our experiments is included along with this paper.
6.2 Statistical analysis
We computed the geometric mean and geometric standard deviations of runtimes, and proportions of optimizer statuses. For the runtime analysis, we only kept instances that were solved to optimality [^3] Sometimes, the optimizer would return a suboptimal status, meaning it cannot prove optimality. If for an instance, our method did converge to optimality, and the direct method returned a suboptimal status, and the objective values were the same up to a tolerance of $10^{-12}$, we consider the suboptimal status to be optimal for the runtime statistical analysis.. We performed paired t-tests on the log of runtimes (for instances where both methods returned optimal) and $\chi^{2}$ tests for statuses.
6.3 Results
Our method consistently outperforms directly solving the optimization problem. This is despite our method involving solving multiple optimization problems. Figure 2 shows that it scales much better with all measures of complexity, and the differences are statistically significant, often at 1% and 0.1% significance.
The difference is particularly striking as models get more complex: in the tree depth experiments, our method is on average 117x faster for trees of depth $6$ (direct mean 82.1 sec, our method mean 0.7 sec), and 391x faster for depth $8$ (direct mean 665.3 sec, our method mean 1.7 sec) – reaching a 1000x speedup on some instances. For neural networks, solving the problem directly times out in every instance, as seen in Figure 3, even for the smallest architecture with 1 hidden layer of 5 neurons. However, our method solves all instances with 1 hidden layer and most instances with 2 hidden layers, although it often times out with 2 hidden layers of 20 neurons each. Figure 3 also shows that the direct method often struggles with proving optimality.
In order to isolate the effects of different parts of our method, we performed an ablation study. We redid the decision trees experiment fixing a tree depth of $8$. We ablated, in turn: (1) finding bounds on $\boldsymbol{\mathbf{\theta}}$, (2) propagating bounds to $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$, (3) initial bounds on $\boldsymbol{\mathbf{v}}$, and (4) tightening bounds on $\boldsymbol{\mathbf{v}}$. Ablating a step means ablating it and all subsequent steps. As shown in Table 1, the biggest contributor to our method’s speedup is obtaining the initial bounds on $\boldsymbol{\mathbf{v}}$, which causes a 153x speedup on average compared to ablating it. This makes sense as it appears in both the bilinear objective and the bilinear constraint.
Table 1: Geometric means of runtimes and number of instances solved for ablations of our method. Ablation numbers refer to ablating: (1) finding bounds on $\boldsymbol{\mathbf{\theta}}$, (2) propagating bounds to $\boldsymbol{\mathbf{\pi}},\boldsymbol{\mathbf{P}},\boldsymbol{\mathbf{r}}$, (3) initial bounds on $\boldsymbol{\mathbf{v}}$, and (4) tightening bounds on $\boldsymbol{\mathbf{v}}$. Lastly, None refers to no ablations. Note that ablating (1) is equivalent to directly solving, and None is equivalent to our full method.
(1) (2) (3) (4) None
Runtime (sec), geom. mean (sd) 710.79 (1.22) 292.36 (1.64) 282.42 (1.59) 1.85 (1.55) 1.89 (1.56)
Instances solved 6/10 8/10 7/10 10/10 10/10
7 Conclusion
We have presented a novel framework for the formal verification of Markov processes with learned parameters, enabling rigorous analysis of systems where transition probabilities and rewards are determined by machine learning models. By formulating the problem as a bilinear program and introducing a novel decomposition and bound propagation scheme, we provide a computationally efficient approach that is orders of magnitude faster than solving the problem directly. Our software, markovml, provides a flexible way of building a Markov process and embedding several common ML models, and optimizing for the three quantities we have discussed in our paper. Our experimental results demonstrate the method’s scalability across varying problem complexities, including state space size, model count, and model complexity. Though our method struggles with increasing neural network size, this is expected as verifying neural networks is computationally difficult. Future work may look at incorporating the latest advances from neural network verification into our framework.
Impact Statement
Our work contributes to improving the safety and reliability of systems that combine machine learning with probabilistic modeling. Such settings arise in diverse fields, including healthcare and engineering. In healthcare, this enables precise analysis of patient risk and cost-effectiveness, supporting more transparent and informed decision-making. For safety-critical autonomous systems, our framework helps ensure that probabilistic failure modes are well understood—even when some components are data-driven—thereby enhancing overall system safety. While our method relies on certain technical assumptions (e.g., MILP-representability of models and feature sets), careful validation in real-world applications can ensure that the formal guarantees are meaningful. Future work should also investigate its applicability to systems of a real-world scale.
Appendix A Formulations of feasibility, reachability, and hitting time
A.1 Feasibility of the total infinite-horizon reward
The total infinite-horizon reward feasibility problem can be written as:
[…]
$\quad W_{\min}\leq\boldsymbol{\mathbf{\pi}}^{\top}\boldsymbol{\mathbf{v}}\leq W_{\max}$
$\quad\boldsymbol{\mathbf{v}}=\lambda\boldsymbol{\mathbf{P}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{r}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\pi}}=\boldsymbol{\mathbf{A_{\pi}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\pi}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{P}})=\boldsymbol{\mathbf{A_{P}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{P}}},$
$\quad\boldsymbol{\mathbf{r}}=\boldsymbol{\mathbf{A_{r}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{r}}},$
$\quad\boldsymbol{\mathbf{C_{\pi}}}\boldsymbol{\mathbf{\pi}}\leq\boldsymbol{\mathbf{d_{\pi}}},$
$\quad\boldsymbol{\mathbf{C_{P}}}\operatorname{vec}(\boldsymbol{\mathbf{P}})\leq\boldsymbol{\mathbf{d_{P}}},$
$\quad\boldsymbol{\mathbf{C_{r}}}\boldsymbol{\mathbf{r}}\leq\boldsymbol{\mathbf{d_{r}}},$
$\quad\sum_{j=1}^{n}P_{ij}=1\quad\forall i\in[n],$
$\quad\sum_{i=1}^{n}\pi_{i}=1,$
$\quad 0\leq P_{ij}\leq 1\quad\forall i,j\in[n]\times[n],$
$\quad 0\leq\pi_{i}\leq 1\quad\forall i\in[n],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
[…]
A.2 Reachability
To reformulate the total reward formulation for reachability to a target state $S\subseteq[n]$ from a set of transient states $T\subseteq[n],T\cap S=\emptyset$, observe that we can simply make the following substitutions: $\boldsymbol{\mathbf{\pi}}$ gets replaced with $\boldsymbol{\mathbf{\tilde{\pi}}}$, which is the initial probability over $T$, $\boldsymbol{\mathbf{P}}$ gets replaced with $\boldsymbol{\mathbf{Q}}$, which is $\boldsymbol{\mathbf{P}}$ restricted to $T$, and $\boldsymbol{\mathbf{r}}$ gets replaced with $\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$, where $\boldsymbol{\mathbf{R}}$ is the transition matrix from $T$ to $S$. Lastly, there is no more discount rate $\lambda$. We also require $\boldsymbol{\mathbf{Q}}$ to be strictly substochastic so that it remains invertible (i.e., the row sums are strictly less than $1$), and $\boldsymbol{\mathbf{R}}$ and $\boldsymbol{\mathbf{\tilde{\pi}}}$ are now substochastic (i.e., the row sums are less than or equal to $1$). The optimization problem is then:
[…]
$\quad\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{Q}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\pi}}=\boldsymbol{\mathbf{A_{\pi}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\pi}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{Q}})=\boldsymbol{\mathbf{A_{Q}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{Q}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{R}})=\boldsymbol{\mathbf{A_{R}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{R}}},$
$\quad\boldsymbol{\mathbf{C_{\tilde{\pi}}}}\boldsymbol{\mathbf{\tilde{\pi}}}\leq\boldsymbol{\mathbf{d_{\tilde{\pi}}}},$
$\quad\boldsymbol{\mathbf{C_{QR}}}[\operatorname{vec}(\boldsymbol{\mathbf{Q}})\operatorname{vec}(\boldsymbol{\mathbf{R}})]\leq\boldsymbol{\mathbf{d_{QR}}},$
$\quad\sum_{j=1}^{|T|}Q_{ij}<1\quad\forall i\in[|T|],$
$\quad\sum_{j=1}^{|S|}R_{ij}\leq 1\quad\forall i\in[|T|],$
$\quad\sum_{i=1}^{|T|}(\tilde{\pi})_{i}\leq 1,$
$\quad 0\leq Q_{ij}<1\quad\forall i,j\in[|T|]\times[|T|],$
$\quad 0\leq R_{ij}\leq 1\quad\forall i,j\in[|T|]\times[|S|],$
$\quad 0\leq(\tilde{\pi})_{i}\leq 1\quad\forall i\in[|T|],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
With a slight abuse of notation, we still maintain the vector $\boldsymbol{\mathbf{v}}$, but is instead now defined as $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$. As well, note that the stochastic constraints have been replaced with substochastic constraints (strict substochastic for $\boldsymbol{\mathbf{Q}}$).
The feasibility problem is:
[…]
$\quad W_{\min}\leq\boldsymbol{\mathbf{\tilde{\pi}}}^{\top}\boldsymbol{\mathbf{v}}\leq W_{\max}$
$\quad\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{Q}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\tilde{\pi}}}=\boldsymbol{\mathbf{A_{\tilde{\pi}}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\tilde{\pi}}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{Q}})=\boldsymbol{\mathbf{A_{Q}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{Q}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{R}})=\boldsymbol{\mathbf{A_{R}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{R}}},$
$\quad\boldsymbol{\mathbf{C_{\tilde{\pi}}}}\boldsymbol{\mathbf{\tilde{\pi}}}\leq\boldsymbol{\mathbf{d_{\tilde{\pi}}}},$
$\quad\boldsymbol{\mathbf{C_{QR}}}[\operatorname{vec}(\boldsymbol{\mathbf{Q}})\operatorname{vec}(\boldsymbol{\mathbf{R}})]\leq\boldsymbol{\mathbf{d_{QR}}},$
$\quad\sum_{j=1}^{|T|}Q_{ij}<1\quad\forall i\in[|T|],$
$\quad\sum_{j=1}^{|S|}R_{ij}\leq 1\quad\forall i\in[|T|],$
$\quad\sum_{i=1}^{|T|}\tilde{\pi}_{i}\leq 1,$
$\quad 0\leq Q_{ij}<1\quad\forall i,j\in[|T|]\times[|T|],$
$\quad 0\leq R_{ij}\leq 1\quad\forall i,j\in[|T|]\times[|S|],$
$\quad 0\leq\tilde{\pi}_{i}\leq 1\quad\forall i\in[|T|],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
A.3 Hitting time
The hitting time to a target set $S\subseteq[n]$ from a set of transient states $T\subseteq[n],T\cap S=\emptyset$ formulation is extremely similar to the reachability formulation, except that we replace $\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$ with $\boldsymbol{\mathbf{1}}$. Hence, $\boldsymbol{\mathbf{R}}$ is no longer a decision variable. Also, now, $\boldsymbol{\mathbf{v}}$ is defined as $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{1}}$. Note again the substochastic constraint on $\boldsymbol{\mathbf{\tilde{\pi}}}$ and the strict substochastic constraint on $\boldsymbol{\mathbf{Q}}$. The optimization problem is:
[…]
$\quad\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{Q}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{1}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\tilde{\pi}}}=\boldsymbol{\mathbf{A_{\tilde{\pi}}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\tilde{\pi}}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{Q}})=\boldsymbol{\mathbf{A_{Q}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{Q}}},$
$\quad\boldsymbol{\mathbf{C_{\tilde{\pi}}}}\boldsymbol{\mathbf{\tilde{\pi}}}\leq\boldsymbol{\mathbf{d_{\tilde{\pi}}}},$
$\quad\boldsymbol{\mathbf{C_{Q}}}\operatorname{vec}(\boldsymbol{\mathbf{Q}})\leq\boldsymbol{\mathbf{d_{Q}}},$
$\quad\sum_{j=1}^{|T|}Q_{ij}<1\quad\forall i\in[|T|],$
$\quad\sum_{i=1}^{|T|}\tilde{\pi}_{i}\leq 1,$
$\quad 0\leq Q_{ij}<1\quad\forall i,j\in[|T|]\times[|T|],$
$\quad 0\leq\tilde{\pi}_{i}\leq 1\quad\forall i\in[|T|],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
The feasibility problem is:
[…]
$\quad W_{\min}\leq\boldsymbol{\mathbf{\tilde{\pi}}}^{\top}\boldsymbol{\mathbf{v}}\leq W_{\max}$
$\quad\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{Q}}\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{1}},$
$\quad\boldsymbol{\mathbf{\theta}}=\begin{bmatrix}\boldsymbol{\mathbf{\theta}}_{1},\boldsymbol{\mathbf{\theta}}_{2},\cdots,\boldsymbol{\mathbf{\theta}}_{k_{f}}\end{bmatrix}^{\top},$
$\quad\boldsymbol{\mathbf{\theta}}_{i}=f_{i}(\boldsymbol{\mathbf{x}})\quad\forall i\in[k_{f}],$
$\quad\boldsymbol{\mathbf{\tilde{\pi}}}=\boldsymbol{\mathbf{A_{\tilde{\pi}}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{\tilde{\pi}}}},$
$\quad\operatorname{vec}(\boldsymbol{\mathbf{Q}})=\boldsymbol{\mathbf{A_{Q}}}\boldsymbol{\mathbf{\theta}}+\boldsymbol{\mathbf{b_{Q}}},$
$\quad\boldsymbol{\mathbf{C_{\tilde{\pi}}}}\boldsymbol{\mathbf{\tilde{\pi}}}\leq\boldsymbol{\mathbf{d_{\tilde{\pi}}}},$
$\quad\boldsymbol{\mathbf{C_{Q}}}\operatorname{vec}(\boldsymbol{\mathbf{Q}})\leq\boldsymbol{\mathbf{d_{Q}}},$
$\quad\sum_{j=1}^{|T|}Q_{ij}<1\quad\forall i\in[|T|],$
$\quad\sum_{i=1}^{|T|}(\pi_{T})_{i}\leq 1,$
$\quad 0\leq Q_{ij}<1\quad\forall i,j\in[|T|]\times[|T|],$
$\quad 0\leq\tilde{\pi}_{i}\leq 1\quad\forall i\in[|T|],$
$\quad\boldsymbol{\mathbf{x}}\in\mathcal{X}.$
Appendix B Omitted Proofs from Section 4
Proof of Lemma 4.1.
For a given component $y_{i}$, we can write the transformation as:
$y_{i}=\sum_{j=1}^{\ell}A_{ij}\theta_{j}+b_{i}$
for all $i=1,2,\ldots,k$.
Each term in the summation depends linearly on $\theta_{j}$. If $A_{ij}\geq 0$, then $A_{ij}\theta_{j}$ is maximized at $\theta_{j}^{\max}$ and minimized at $\theta_{j}^{\min}$. Vice versa, if $A_{ij}<0$, then $A_{ij}\theta_{j}$ is maximized at $\theta_{j}^{\min}$ and minimized at $\theta_{j}^{\max}$.
Applying this to all terms in the summation, we get the required bounds. ∎
Proof of Lemma 4.2.
Note that the matrix $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})$ is a (non-singular) M-matrix. By Theorem 2.3 in Chapter 6 of (Berman & Plemmons, 1994), each element of the matrix is non-negative, so is $\geq 0$.
Next, consider the row sum of $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}$ for a row $i$: $\sum_{j}(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}_{ij}$, and then expand into the Neumann series: $\sum_{j}\sum_{k=0}^{\infty}\lambda^{k}(\boldsymbol{\mathbf{P}}^{k})_{ij}$. We can exchange the order of summation as all terms are non-negative, and then we have our sum is $\sum_{k=0}^{\infty}\lambda^{k}\sum_{j}(\boldsymbol{\mathbf{P}}^{k})_{ij}$. Every positive integer exponent of a row-stochastic matrix is also row-stochastic, so $\sum_{j}(\boldsymbol{\mathbf{P}}^{k})_{ij}=1$, for all $k\in\mathbb{N}$. Therefore, this simplifes to the geometric series $\sum_{k=0}^{\infty}\lambda^{k}$, which equals $\frac{1}{1-\lambda}$. As each row sum is $\frac{1}{1-\lambda}$, and each element is non-negative, each element must also be at most $\frac{1}{1-\lambda}$. ∎
Proof of Lemma 4.3.
Recall that $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}\boldsymbol{\mathbf{r}}$, so $v_{i}=\sum_{j=1}^{n}(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}_{ij}r_{j}$. Thus, we can see $v_{i}$ is a positive linear combination of the elements of $\boldsymbol{\mathbf{r}}$, where the nonnegative weights are from $(\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}})^{-1}$ and sum to $1/(1-\lambda)$. To maximize this combination, we assign all the weight to the maximum upper bound of $\boldsymbol{\mathbf{r}}$, and to minimize it, we assign all the weight to the minimum lower bound of $\boldsymbol{\mathbf{r}}$. Hence, we have the required bounds. ∎
Proof of Theorem 4.4.
Let $\boldsymbol{\mathbf{M}}=\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}}$. Clearly, $\boldsymbol{\mathbf{M}}$ is bounded element-wise by $\boldsymbol{\mathbf{M^{\max}}}=\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P^{\min}}}$ and $\boldsymbol{\mathbf{M^{\min}}}=\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P^{\max}}}$. The upper bound $\boldsymbol{\mathbf{M^{\max}}}$ has no positive off-diagonal elements, as all of the elements of $\boldsymbol{\mathbf{P^{\max}}}$ are between 00 and $1$. So, the only thing we need to check is that $\boldsymbol{\mathbf{M^{\min}}}$ is an M-matrix.
From the definition of an M-matrix (Berman & Plemmons, 1994), $\boldsymbol{\mathbf{M^{\min}}}$ is an M-matrix if and only if $1\geq\rho(\lambda\boldsymbol{\mathbf{P^{\max}}})=\lambda\rho(\boldsymbol{\mathbf{P^{\max}}})$. This completes the proof. ∎
Appendix C Gauss-Seidel Method for Interval Matrices
Our exposition here follows section 5.7 of Horáček (2019). Suppose we have a linear system $\boldsymbol{\mathbf{A}}\boldsymbol{\mathbf{x}}=\boldsymbol{\mathbf{b}}$, where $\boldsymbol{\mathbf{A}}$ is an interval matrix and $\boldsymbol{\mathbf{b}}$ is an interval vector. The set of $\boldsymbol{\mathbf{x}}$ that satisfies this system is referred to as the solution set. It is in general complicated, so we hope to find a rectangular enclosure of the solution set. The smallest such enclosure is referred to as the hull of the solution set.
Suppose we have an initial enclosure of the solution set, $\boldsymbol{\mathbf{x}}^{(0)}$. In general this takes some effort to obtain, but we obtained this for our specific problems. Then, the interval Gauss-Seidel method proceeds as follows:
  1. 1.
    For each variable $x_{i}$, compute a new enclosure of the solution set as:
    $y_{i}^{(k+1)}=\frac{1}{A_{ii}}\left(b_{i}-\sum_{j<i}A_{ij}x_{j}^{(k+1)}-\sum_{j>i}A_{ij}x_{j}^{(k)}\right)$
  2. 2.
    Update the enclosure of the solution set: $\boldsymbol{\mathbf{x}}^{(k+1)}=\boldsymbol{\mathbf{x}}^{(k)}\cap\boldsymbol{\mathbf{y}}^{(k+1)}$.
  3. 3.
    Repeat steps 1 and 2 until stopping criterion is met (either a maximum number of iterations or the difference in subsequent enclosures is less than some tolerance).
All operations above are interval arithmetic operations. For real intervals, we have: $[a,b]+[c,d]=[a+c,b+d]$, $[a,b]-[c,d]=[a-d,b-c]$, $[a,b]\cdot[c,d]=[\min(ac,ad,bc,bd),\max(ac,ad,bc,bd)]$. Division is defined as multiplication by the reciprocal, i.e., $[a,b]/[c,d]=[a,b]\cdot[1/d,1/c]=[\min(a/c,a/d,b/c,b/d),\max(a/c,a/d,b/c,b/d)]$, if $0\notin[c,d]$.
Note that the zero division issue is not a concern for us, as the diagonal elements of $\boldsymbol{\mathbf{I}}-\lambda\boldsymbol{\mathbf{P}}$, in the total reward setting, have lower bounds always strictly positive, due to the discount factor $\lambda\in(0,1)$. In the reachability and hitting time settings, the diagonal elements of $\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}}$ have lower bounds always strictly positive, due to the strict substochastic property of $\boldsymbol{\mathbf{Q}}$. In practice, we use a small numerical offset to handle this.
Appendix D Solving other formulations
D.1 Reachability and hitting time
The reachability and hitting time formulations are in Appendix A. The key difference that affects our theoretical results and algorithms is that the matrix $\boldsymbol{\mathbf{Q}}$, which is a submatrix of $\boldsymbol{\mathbf{P}}$ restricted to the transient states, is strictly substochastic, and we don’t have a discount factor $\lambda$ anymore. This introduces some technical difficulties in guaranteeing invertibility of $\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}}$. However, all our theoretical results all follow through with minor modifications.
Lemma D.1.
For a strictly row-substochastic matrix $\boldsymbol{\mathbf{Q}}$ with all elements in $[0,1]$:
  1. 1.
    The spectral radius of $\boldsymbol{\mathbf{Q}}$ is strictly less than $1$.
  2. 2.
    The matrix $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})$ is a (non-singular) M-matrix.
  3. 3.
    The matrix $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$ has all non-negative elements.
  4. 4.
    The row sums of $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$ are at most $1/(1-\alpha)$, where $\alpha=\max_{i}\sum_{j}Q_{ij}$ is the maximum row sum of $\boldsymbol{\mathbf{Q}}$.
  5. 5.
    Each element of $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$ is in the interval $[0,1/(1-\alpha)]$.
Proof.
The row sum of $\boldsymbol{\mathbf{Q}}$ is strictly less than $1$, so, by a well-known result in linear algebra, the spectral radius of $\boldsymbol{\mathbf{Q}}$ is strictly less than $1$. It follows that $\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}}$ is a (non-singular) M-matrix by the definition of an M-matrix (Berman & Plemmons, 1994), and hence that the inverse has all non-negative elements by Theorem 2.3 in Chapter 6 of Berman & Plemmons (1994). ∎
Note that the maximum row sum $\alpha$ is strictly less than $1$, so the formula is well-defined. We use this result to obtain the appropriate initial bounds on $\boldsymbol{\mathbf{v}}$ for the reachability and hitting time formulations.
Corollary D.2.
Given a Markov reachability problem with element-wise bounds on $\boldsymbol{\mathbf{Q}}$ and $\boldsymbol{\mathbf{R}}$, as defined in Appendix A, the initial bounds on $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$ are:
$0\leq v_{i}\leq\frac{1}{1-\alpha}\max_{j}\sum_{k}R_{jk}^{\max}$
where $\alpha=\max_{i}\sum_{j}Q_{ij}^{\max}$ is the maximum row sum of $\boldsymbol{\mathbf{Q}}$.
Proof.
First, note that the vector $\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$ has elements which are the row sums of $\boldsymbol{\mathbf{R}}$. So, for each $i$, $\sum_{j}R_{ij}^{\min}\leq(\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}})_{i}\leq\sum_{j}R_{ij}^{\max}$. Next, noting that the row sums of $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$ are at most $1/(1-\alpha)$, we have, by a similar argument as in the proof of Lemma 4.3, that the upper bound on $v_{i}$ is:
$v_{i}\leq\frac{1}{1-\alpha}\max_{j}\sum_{k}R_{jk}^{\max}$
For the lower bound, unlike the argument in Lemma 4.3, we don’t know what the row sums of $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$ actually are, so we use the smallest possible row sum, which is 00. So, we have:
$v_{i}\geq 0$
Corollary D.3.
Given a Markov hitting time problem with element-wise bounds on $\boldsymbol{\mathbf{Q}}$, as defined in Appendix A, the initial bounds on $\boldsymbol{\mathbf{v}}=(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}\boldsymbol{\mathbf{1}}$ are:
$0\leq v_{i}\leq\frac{1}{1-\alpha}$
where $\alpha=\max_{i}\sum_{j}Q_{ij}^{\max}$ is the maximum row sum of $\boldsymbol{\mathbf{Q}}$.
Proof.
The bound simply follows from the fact that $\boldsymbol{\mathbf{v}}$ is the row sums of $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})^{-1}$. ∎
Note, however, that we do not know what the row sums of $\boldsymbol{\mathbf{Q}}$ are, because it is not fixed. However, in practice, in order to assert the strict row-substochastic property, we use a small numerical offset, which is what we use in our implementation. Specifically, we set the row sums of $\boldsymbol{\mathbf{Q}}$ to be at most $1-\epsilon$, where $\epsilon$ is by default set to $10^{-6}$ in our software implementation. Based on this, we can set $\alpha$ to be $1-\epsilon$.
Next, we can tighten the bounds on $\boldsymbol{\mathbf{v}}$ by forming the appropiate linear systems: $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{R}}\boldsymbol{\mathbf{1}}$ for reachability and $(\boldsymbol{\mathbf{I}}-\boldsymbol{\mathbf{Q}})\boldsymbol{\mathbf{v}}=\boldsymbol{\mathbf{1}}$ for hitting time. In both cases, though $\boldsymbol{\mathbf{Q}}$ is a submatrix of $\boldsymbol{\mathbf{P}}$, it is square, so we can analyze the similar interval M-matrix condition for the Gauss-Seidel method to achieve the hull of the solution set.
Corollary D.4.
Let $\boldsymbol{\mathbf{Q}}$ be a row-substochastic matrix bounded element-wise by $\boldsymbol{\mathbf{Q^{\min}}}$ and $\boldsymbol{\mathbf{Q^{\max}}}$. Then, it is an interval M-matrix if and only if $\rho(\boldsymbol{\mathbf{Q^{\max}}})\leq 1$.
Proof.
The proof is identical to that of Theorem 4.4, but without the discount factor. ∎
Based on the above extensions of our theoretical results, we now have similar algorithms for reachability and hitting time. As usual, we first optimize for bounds on $\boldsymbol{\mathbf{\theta}}$, then we propagate to $\boldsymbol{\mathbf{Q}}$ and $\boldsymbol{\mathbf{R}}$ (if applicable), obtain initial bounds on $\boldsymbol{\mathbf{v}}$ as above, and then apply the Gauss-Seidel method to tighten the bounds on $\boldsymbol{\mathbf{v}}$.
D.2 Feasibility
Our method easily extends to the feasibility version of all three problems. We still obtain all the necessary bounds as in the optimization version. At the end, when we would optimize for the quantity of interest, we simply check if the overall problem is feasible, given the bounds on the quantity of interest.
Appendix E Implementation Details
We release markovml along with this paper, with full documentation and tutorials. We also include the code to reproduce our experiments. Below we discuss some important features.
E.1 Supported Models
Currently, the implementation supports the following models from sklearn:
  • Linear regression (sklearn.linear_model.LinearRegression)
  • Ridge regression (sklearn.linear_model.Ridge)
  • Lasso regression (sklearn.linear_model.Lasso)
  • Logistic regression (sklearn.linear_model.LogisticRegression)
  • Decision tree regression (sklearn.tree.DecisionTreeRegressor)
  • Decision tree classifier (sklearn.tree.DecisionTreeClassifier)
  • Random forest regression (sklearn.ensemble.RandomForestRegressor)
  • Random forest classifier (sklearn.ensemble.RandomForestClassifier)
  • Multi-layer perceptron regression (sklearn.neural_network.MLPRegressor)
  • Multi-layer perceptron classifier (sklearn.neural_network.MLPClassifier)
From pytorch, we support neural networks build as pytorch.nn.Sequential models with ReLu or linear layers, with possibly a softmax layer at the end for classifiers.
Lastly, we implemented a model called DecisionRules, which allows the user to specify a series of “if-then” rules specified in natural language, e.g., “if age ¿= 65 then 0.8”. This enables encoding tables from the literature, like age-stratified mortalities. Our code parses these strings into an internal representation.
E.2 MILP formulations
We make use of the MILP formulations from gurobi-machinelearning. At time of writing, gurobi-machinelearning did not support the softmax function, although it did support the logistic function. Hence, for multilayer perceptron classifiers from sklearn and for neural network classifers from torch, we implemented the MILP formulation ourselves. We also implemented the MILP formulation of our DecisionRules model through a series of logical implications.
E.3 Code Example
In only a few lines, a user can build a Markov process, integrate an ML model, and optimize the reward.
1from markovml.markovml import MarkovReward
2from sklearn.linear_model import LogisticRegression
3import numpy as np
4
5# Create a Markov process with 2 states and 2 features
6mrp = MarkovReward(n_states=2, n_features=2)
7
8# fix some of the parameters
9mrp.set_r([1, 0])
10mrp.set_pi([1, 0])
11
12# train a classifier
13X = np.random.rand(100, 2)
14y = np.random.randint(0, 2, 100)
15clf = LogisticRegression().fit(X, y)
16
17# add the classifier to the Markov process
18mrp.add_ml_model(clf)
19
20# link it to the transition probabilities
21# this means the first output of the first ML model is the probability of transitioning to the second state
22mrp.set_P([[1 - mrp.ml_outputs[0][0], mrp.ml_outputs[0][0]], [0, 1]])
23
24# set bounds on the features
25mrp.add_feature_constraint(mrp.features[0] >= 65)
26mrp.add_feature_constraint(mrp.features[1] >= 100)
27
28# optimize the reward
29mrp.optimize(sense="max")
Listing 1: Optimizing the total reward of a Markov process with parameters given by a logistic regression model
All the necessary steps, like checking validity of values, adding the MILP formulation of the ML model to the optimization problem, as well as running our decomposition and bound propagation scheme, are handled in the background by our package.